OpenStudy (nightmarenight):

A cylindrical container has a volume of 36 cm³. Another container made out of the same material is similar to the first container with a height that is three-fourths the height of the first container. To the nearest tenth of a cm³, what is the volume of the second container?

OpenStudy (anonymous):

trick question ?=27

OpenStudy (anonymous):

Formula for volume of a cylinder, \[V=\pi r^2h\] Now for cylinder 1, \[V_{1}=36\]\[r_{1}=r\](let) \[h_{1}=h\] (let) Cylinder 2, \[V_{2}=x\](let)\[r_{2}=r\](same radius for both cylinders) \[h_{2}=\frac{3}{4}h\](this is given in question) Now the next step is to make the equations, Cylinder 1, \[V_{1}=\pi r_{1}^2h_{1}\]\[36=\pi r^{2}h\] Cylinder 2, \[V_{2}=\pi r_{2}^2h_{2}\]\[x=\frac{3}{4}\pi r^{2}h\] Now you have 2 equations and 2 variables these can be solved using elimination or substitution methods

OpenStudy (anonymous):

A shorter method, A cylinder's volume is directly proportional to it's height Therefore we can write \[V=kh\] where k is your constant of proportionality Cylinder 1, \[V_{1}=kh_{1}\] Cylinder 2, \[V_{2}=kh_{2}\] Divide equation 2 by equation 1, \[\frac{V_{2}}{V_{1}}=\frac{k}{k} \times \frac{h_{2}}{h_{1}}\]\[\frac{V_{2}}{V_{1}}=\frac{h_{2}}{h_{1}}\]\[V_{2}=V_{1}.\frac{h_{2}}{h_{1}}\] use \[h_{2}=\frac{3}{4}h_{1}\] and\[V_{1}=36\] In short since the radius is constant and volume is directly proportional to height, if height is 3/4th then volume would also become 3/4th